# LLM-SmartAudit: Advanced Smart Contract Vulnerability Detection

### v1

@@ Abstract
I Introduction
@@
Recent advances in smart contracts have marked significant progress in areas such as security, finance, and governance. These are self-executing contracts with terms mutually agreed upon by participants, enacted through predefined actions. However, the immutable nature of blockchain technology inherently makes smart contracts more susceptible to software attacks. Over the past years, there have been numerous high-profile vulnerabilities and exploits in smart contracts, such as the DAO attack [1]. Consequently, the development of secure smart contracts continues to pose significant challenges.
@@
Research in Large Language Models (LLMs) has made significant advancements in fields such as Natural Language Processing [2], Computer Vision [3], Code Generation [4], and various AI tasks [5, 6]. A survey [7] indicates a significant increase in LLMs adoption over the past five years, accompanied by a rapid rise in software engineering. LLM tools are primarily categorized into commercial products and open-source initiatives. State-of-the-art (SOTA) commercial LLMs include GPT [8], Claude [9], and Gemini [10] models, while prominent open-source projects feature Llama [11], Mixtral [12], and GPT-NeoX [13]. Both commercial and open-source LLMs demonstrate significant potential in handling complex tasks. GPT, one of the pioneering commercial LLMs, has exhibited remarkable proficiency in natural language understanding, context processing, and code comprehension, outperforming numerous traditional methods [14].
@@
LLMs, trained on vast text data, excel at predicting likely token sequences based on input. Unlike traditional systems with fixed rules or vulnerability databases, LLMs generate probabilistic outputs. While powerful, this approach can sometimes lead to inaccuracies in vulnerability identification [15, 16, 17]. This paper investigate the question: ‘Can multi-agent conversations enhance LLMs’ capabilities in detecting smart contract vulnerabilities?’.
@@
We propose a virtual chat-powered smart contract audit framework utilizing a multi-agent conversation system. This approach is based on three key rationales. Firstly, LLMs’ ability to incorporate feedback and engage in cooperative conversations enables a dynamic, interactive, and iterative approach to identifying and addressing smart contract vulnerabilities. Secondly, the collaboration of agents with specialized knowledge and analytical skills enhances the factual accuracy and reasoning precision of the audit process [18]. Lastly, multi-agent interaction mitigates the ‘degeneration-of-thought’ issue common in single-agent self-reflection processes [19]. This approach allows agents to challenge and complement each other’s viewpoints, resulting in a more balanced and comprehensive analysis. Supporting this, Wu et al. [20] demonstrated that multi-agent systems like AutoGen and CAMEL [21] outperform single-agent systems such as AutoGPT [22] and LangChain Agents [23] across various applications. These findings underscore the potential effectiveness of multi-agent systems in conducting smart contract audits.
@@
  • The Power of Multiple LLM Agents: Our research reveals that each LLM Agent, each assembled with the specific capabilities and roles, specialize in distinct areas of the security auditing, including contract code analysis, vulnerability identification, and security summary. These specialized agents, guided by step-by-step instructions, perform in-depth analysis within their respective domains, yielding more accurate and comprehensive results. Furthermore, LLM Agents collaborate seamlessly, exchanging data and insights to provide a holistic view of the security landscape.
@@
The rest of the paper is organized as follows. Section II presents the current challenges in smart contract analysis and surveys the landscape of existing tools. Section III details the architecture and operational mechanism of LLM-SmartAudit. Section IV presents a comprehensive evaluation of LLM-SmartAudit, detailing the standardized datasets, evaluation criteria, experimental results, and related works. Section V discusses the findings and addresses potential threats to validity. Finally, Section VI concludes the paper, summarizing the contributions and offering insights into the future of smart contract analysis tools.
II Backgrounds
II-A Smart Contract Security
@@
Smart contracts, self-executing agreements encoded in software, facilitate the management and execution of digital assets within blockchain networks [24]. These contracts not only define rules and penalties in a similar way to traditional contracts but also automatically enforce these obligations. Although Nick Szabo first proposed the concept in 1994 [25], smart contracts become practically implementable with Ethereum’s launch in 2015. A recent survey [26] reveals a rapid increase in the number of Solidity contracts over the past five years. This growth reflects the expanding range of smart contract applications across sectors such as decentralized finance (DeFi) [27], insurance, and lending platforms. Notably, the DeFi sector has experienced significant growth, with its peak total value locked (TVL) reaching 179 billion USD on November 9, 2021. Of this, Ethereum accounts for 108.9 billion USD, representing 60% of the total DeFi TVL [28].
@@
The substantial asset value managed by smart contracts underscores their critical security importance. However, a defining characteristic of Solidity smart contracts is their post-deployment immutability on the Ethereum network, presenting significant security management challenges. Unlike traditional software, where patches or updates can rectify bugs or flaws, smart contracts lack this flexibility. Consequently, vulnerabilities discovered post-deployment remain unfixable in the existing contract, potentially leading to substantial financial losses if exploited by malicious actors. According to Zhou et al. [26], smart contracts have been the target of numerous high-profile attacks, resulting in losses exceeding 3.24 billion USD from April 2018 to April 2022.
II-B Automated Security Analysis
@@
With the rise in security incidents and high-profile attacks, diverse smart contract analysis tools have been developed. These tools are designed to systematically detect vulnerabilities, enforce best practices, and identify potential security risks inherent in smart contracts. By leveraging these tools, developers can proactively mitigate issues before malicious exploitation, significantly reducing security breach risks and ensuring contract execution integrity. These tools employ advanced techniques such as formal verification, symbolic execution, intermediate representation (IR), and machine learning to enhance their effectiveness [29, 30].
@@
Despite these advancements, substantial challenges persist in smart contract security analysis. A primary concern is the complexity and diversity of vulnerabilities, making it difficult for any single tool to be universally effective. Each tool has its own strengths and limitations. For instance, tools relying on formal verification excel at ensuring contracts adhere to specified requirements but may fall short in detecting security flaws like reentrancy or gas limit issues. Complex logic vulnerabilities still necessitate human auditors [31], introducing additional challenges. However, the fees charged by traditional smart contract auditing firms are prohibitively high. Basic audits from firms like CertiK start at 500 USD, while more reputable companies such as Trail of Bits charge between 5,000 USD to 15,000 USD as a starting point [32].
II-C LLMs for Vulnerability Prediction
@@
LLMs, such as GPT and Claude, are a specific type of generative AI focused on text generation [33]. These models are termed "large" due to their vast number of parameters, enabling them to comprehend and produce human language with remarkable coherence and contextual appropriateness. Pre-trained on diverse internet-based text sources, they can produce text that often mirrors the quality and style of human writing. LLMs have demonstrated the ability to grasp grammatical structures, word meanings, and basic logical reasoning in human languages.
@@
LLMs have demonstrated excellence in specific downstream tasks, including code completion, code analysis, and vulnerability detection. By leveraging their code comprehension and generation capabilities, these models can identify vulnerabilities, verify compliance, and assess logical correctness. Their effectiveness is further enhanced through advanced prompting techniques like chain-of-thought (CoT) or few-shot. Chen et al. [17] have proven that LLMs (GPT-2, T5), trained with a high-quality dataset consisting of 18,945 vulnerable C/C++ functions, outperform other machine learning methods, including Graph Neural Networks, in vulnerability prediction. Khare et al. [34] found that proper prompting strategies that involve step-by-step analysis significantly improve the performance of LLMs (GPT, CodeLlama) in detecting security vulnerabilities in programming languages such as Java and C/C++.
III LLM-SmartAudit System
@@
Figure 1: Muliti-agent Conversation Framework
III-A Framework
@@
III-A1 Role Specialization
The framework adopts a role-playing methodology to define and specialize the function of each agent within the audit process. By utilizing the inception prompting technique [21], the system enables agents to effectively assume and fulfill their designated roles. Agents are assigned specific capabilities and roles, either through repurposing existing agents or extending their functionalities. These roles encompass sending messages and receiving information from other agents, facilitating the initiation and continuation of inter-agent conversations.
The framework deploys a combination of an LLM-powered assistant agent and a user agent in an assistant-user cooperative scenario. The assistant agent, powered by LLMs [8], generates solutions and communicates these to the user agent. The user agent then executes the assistant’s recommendations, providing feedback to the assistant agent.
@@ III-A1 Role Specialization
III-A2 Action Execution
After role specialization, assistant and user agents collaborate in an instruction-following manner to complete the assigned tasks. Human users initiate the action execution by inputting contract codes into the system. The system segments this code into a task queue, which is systematically processed through three distinct phases: Contract Analysis, Vulnerability Identification, and Security Summary. The Contract Analysis phase involves preliminary assessments of the contract’s purpose and structure. During Vulnerability Identification phase, agents collaboratively identify and describe potential security weaknesses. Finally, the Security Summary phase generates a comprehensive audit report.
@@ III-A2 Action Execution
(a) Design of Zero-shot-CoT Prompting
(b) Design of Few-shot-CoT Prompting
Figure 2: Enhanced Prompting Strategies
III-B Task Queue
@@
III-B1 Inception Prompt
@@
Prompt engineering is another critical component, instrumental in achieving role specialization and effective task execution. To facilitate this, the system employs inception prompting [21], a technique that describes the tasks and responsibilities of the assistant and user agents within the framework. An inception prompt comprises three key components: specified task prompt, assistant agent prompt, and user agent prompt. This structure enables the system to transform preliminary ideas into specific, actionable tasks, leveraging the specialized capabilities of each agent. Building on this foundation, the framework incorporates two advanced prompting techniques tailored for the complexities of smart contract analysis: zero-shot-CoT (Chain of Thought) and few-shot-CoT prompting.
Zero-shot-CoT prompting leverages the model’s ability to generate responses to new prompts with encouraging sequential, logical processing. Zero-shot prompting allows models to respond to unfamiliar prompts, but often results in surface-level analyses [35]. This limitation is particularly problematic for complex tasks requiring detailed understanding and logical thought progression. To address this, LLM-SmartAudit adopts the zero-shot-CoT approach [36], which appends the phrase “Let’s think step by step" to the original prompt.
Figure 2(a) showcases an example of the zero-shot-CoT prompt template, tailored for broad-spectrum analysis. The specified task prompt clearly outlines the primary objective for the assistant agents, which is to identify all types of vulnerabilities within the scope of GPT’s knowledge base. The prompt also specifies the expected output format, including vulnerability type and detailed description, while incorporating constraints to mitigate potential undesirable behaviors. The assistant agent prompt and user agent prompt are well structured, each providing specific information about assigned roles and tasks. These prompts are strategically designed to engage two specialized agents, leveraging their capabilities to collaboratively identify vulnerabilities in smart contract.
Few-shot-CoT prompting enhances the model’s task comprehension by providing context-specific examples, particularly crucial for complex domains. While zero-shot is effective for many tasks, it often falls short in complex domains [11]. LLMs often lack domain-specific effectiveness in zero-shot scenarios, limiting their performance in specialized fields. Few-shot addresses this limitation by providing context-specific examples, facilitating in-context learning and enhancing the model’s task comprehension. However, conventional few-shot prompting may not inherently guide LLMs to engage in deep, step-by-step reasoning necessary for complex tasks. To address this, LLM-SmartAudit implements the few-shot-CoT approach. This method combines CoT reasoning with few-shot examples, prompting LLMs to both consider the provided context and methodically analyze the problem.
@@
Figure 2(b) presents an example of the specified task prompt for few-shot-CoT, tailored for focused and detailed analysis of smart contract. This prompt establishes a clear objective for the agents, concentrating on identifying Transactions Order Dependence (TOD) vulnerabilities. It directs agents to examine the contract’s logic and critical functions, particularly those involving fund transfers, resource allocation, or gas price manipulation, for TOD vulnerabilities. If a TOD vulnerability is identified, the assistant agent must provide a detailed description of the vulnerability and its potential impact. Conversely, if no TOD vulnerability is found, the agent must output “NO TOD".
III-B2 Execution Mode

### v2

@@ Abstract.
datasets, neural networks, gaze detection, text tagging
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; October 03–05, 2024; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computer systems organization Embedded systemsccs: Computer systems organization Redundancyccs: Computer systems organization Roboticsccs: Networks Network reliability
1. Introduction
@@
Recent advances in smart contracts have marked significant progress in areas such as security, finance, and governance. These are self-executing contracts with terms mutually agreed upon by participants, enacted through predefined actions. However, the immutable nature of blockchain technology inherently makes smart contracts more susceptible to software attacks. Over the past years, there have been numerous high-profile vulnerabilities and exploits in smart contracts, such as the DAO attack (Atzei et al., 2017). Consequently, the development of secure smart contracts continues to pose significant challenges.
@@
Research in Large Language Models (LLMs) has made significant advancements in fields such as Natural Language Processing (Danilevsky et al., 2020), Computer Vision (Ramesh et al., 2022), Code Generation (Dong et al., 2023), and various AI tasks (Shen et al., 2024; Qian et al., 2023). A survey (Zhang et al., 2023a) indicates a significant increase in LLMs adoption over the past five years, accompanied by a rapid rise in software engineering. LLM tools are primarily categorized into commercial products and open-source initiatives. State-of-the-art (SOTA) commercial LLMs include GPT (OpenAI, 2022), Claude (cla, 2024), and Gemini (Team et al., 2024) models, while prominent open-source projects feature Llama (Touvron et al., 2023), Mixtral (Jiang et al., 2024), and GPT-NeoX (Black et al., 2022). Both commercial and open-source LLMs demonstrate significant potential in handling complex tasks. GPT, one of the pioneering commercial LLMs, has exhibited remarkable proficiency in natural language understanding, context processing, and code comprehension, outperforming numerous traditional methods (Xu et al., 2022).
@@
LLMs, trained on vast text data, excel at predicting likely token sequences based on input. Unlike traditional systems with fixed rules or vulnerability databases, LLMs generate probabilistic outputs. While powerful, this approach can sometimes lead to inaccuracies in vulnerability identification (Azamfirei et al., 2023; Kang et al., 2023; Chen et al., 2023a). This paper investigate the question: ‘Can multi-agent conversations enhance LLMs’ capabilities in detecting smart contract vulnerabilities?’.
@@
We propose a virtual chat-powered smart contract audit framework utilizing a multi-agent conversation system. This approach is based on three key rationales. Firstly, LLMs’ ability to incorporate feedback and engage in cooperative conversations enables a dynamic, interactive, and iterative approach to identifying and addressing smart contract vulnerabilities. Secondly, the collaboration of agents with specialized knowledge and analytical skills enhances the factual accuracy and reasoning precision of the audit process (Salewski et al., 2023). Lastly, multi-agent interaction mitigates the ‘degeneration-of-thought’ issue common in single-agent self-reflection processes (Liang et al., 2023). This approach allows agents to challenge and complement each other’s viewpoints, resulting in a more balanced and comprehensive analysis. Supporting this, Wu et al. (Wu et al., 2023) demonstrated that multi-agent systems like AutoGen and CAMEL (Li et al., 2023) outperform single-agent systems such as AutoGPT (Yang et al., 2023) and LangChain Agents (LangChain, 2023) across various applications. These findings underscore the potential effectiveness of multi-agent systems in conducting smart contract audits.
@@
  • The Power of Multiple LLM Agents: Our research reveals that each LLM Agent, each assembled with the specific capabilities and roles, specialize in distinct areas of the security auditing, including contract code analysis, vulnerability identification, and Comprehresive Report. These specialized agents, guided by step-by-step instructions, perform in-depth analysis within their respective domains, yielding more accurate and comprehensive results. Furthermore, LLM Agents collaborate seamlessly, exchanging data and insights to provide a holistic view of the security landscape.
@@
The rest of the paper is organized as follows. Section 2 presents the current challenges in smart contract analysis and surveys the landscape of existing tools. Section 3 details the architecture and operational mechanism of LLM-SmartAudit. Section 4 presents a comprehensive evaluation of LLM-SmartAudit, detailing the standardized datasets, evaluation criteria, experimental results, and related works. Section 5 discusses the findings and addresses potential threats to validity. Finally, Section 6 concludes the paper, summarizing the contributions and offering insights into the future of smart contract analysis tools.
2. Backgrounds
2.1. Smart Contract Security
@@
Smart contracts, self-executing agreements encoded in software, facilitate the management and execution of digital assets within blockchain networks (Luu et al., 2016). These contracts not only define rules and penalties in a similar way to traditional contracts but also automatically enforce these obligations. Although Nick Szabo first proposed the concept in 1994 (Szabo, 1996), smart contracts become practically implementable with Ethereum’s launch in 2015. A recent survey (Zhou et al., 2023) reveals a rapid increase in the number of Solidity contracts over the past five years. This growth reflects the expanding range of smart contract applications across sectors such as decentralized finance (DeFi) (Berg et al., 2022), insurance, and lending platforms. Notably, the DeFi sector has experienced significant growth, with its peak total value locked (TVL) reaching 179 billion USD on November 9, 2021. Of this, Ethereum accounts for 108.9 billion USD, representing 60% of the total DeFi TVL (DeFillama, 2023).
@@
The substantial asset value managed by smart contracts underscores their critical security importance. However, a defining characteristic of Solidity smart contracts is their post-deployment immutability on the Ethereum network, presenting significant security management challenges. Unlike traditional software, where patches or updates can rectify bugs or flaws, smart contracts lack this flexibility. Consequently, vulnerabilities discovered post-deployment remain unfixable in the existing contract, potentially leading to substantial financial losses if exploited by malicious actors. According to Zhou et al. (Zhou et al., 2023), smart contracts have been the target of numerous high-profile attacks, resulting in losses exceeding 3.24 billion USD from April 2018 to April 2022.
2.2. Automated Security Analysis
@@
With the rise in security incidents and high-profile attacks, diverse smart contract analysis tools have been developed. These tools are designed to systematically detect vulnerabilities, enforce best practices, and identify potential security risks inherent in smart contracts. By leveraging these tools, developers can proactively mitigate issues before malicious exploitation, significantly reducing security breach risks and ensuring contract execution integrity. These tools employ advanced techniques such as formal verification, symbolic execution, intermediate representation (IR), and machine learning to enhance their effectiveness (Tolmach et al., 2021; Chen et al., 2020).
@@
Despite these advancements, substantial challenges persist in smart contract security analysis. A primary concern is the complexity and diversity of vulnerabilities, making it difficult for any single tool to be universally effective. Each tool has its own strengths and limitations. For instance, tools relying on formal verification excel at ensuring contracts adhere to specified requirements but may fall short in detecting security flaws like reentrancy or gas limit issues. Complex logic vulnerabilities still necessitate human auditors (cod, 2023), introducing additional challenges. However, the fees charged by traditional smart contract auditing firms are prohibitively high. Basic audits from firms like CertiK start at 500 USD, while more reputable companies such as Trail of Bits charge between 5,000 USD to 15,000 USD as a starting point (Mateusz, 2024).
2.3. LLMs for Vulnerability Prediction
@@
LLMs, such as GPT and Claude, are a specific type of generative AI focused on text generation (Epstein et al., 2023). These models are termed ”large” due to their vast number of parameters, enabling them to comprehend and produce human language with remarkable coherence and contextual appropriateness. Pre-trained on diverse internet-based text sources, they can produce text that often mirrors the quality and style of human writing. LLMs have demonstrated the ability to grasp grammatical structures, word meanings, and basic logical reasoning in human languages.
@@
LLMs have demonstrated excellence in specific downstream tasks, including code completion, code analysis, and vulnerability detection. By leveraging their code comprehension and generation capabilities, these models can identify vulnerabilities, verify compliance, and assess logical correctness. Their effectiveness is further enhanced through advanced prompting techniques like chain-of-thought (CoT) or few-shot. Chen et al. (Chen et al., 2023a) have proven that LLMs (GPT-2, T5), trained with a high-quality dataset consisting of 18,945 vulnerable C/C++ functions, outperform other machine learning methods, including Graph Neural Networks, in vulnerability prediction. Khare et al. (Khare et al., 2023) found that proper prompting strategies that involve step-by-step analysis significantly improve the performance of LLMs (GPT, CodeLlama) in detecting security vulnerabilities in programming languages such as Java and C/C++.
Figure 1. Muliti-agent Conversation Framework
3. LLM-SmartAudit System
@@
3.1. Framework
@@
3.1.1. Role Specialization
The framework adopts a role-playing methodology to define and specialize the function of each agent within the audit process. By utilizing the inception prompting technique (Li et al., 2023), the system enables agents to effectively assume and fulfill their designated roles. Agents are assigned specific capabilities and roles, either through repurposing existing agents or extending their functionalities. These roles encompass sending messages and receiving information from other agents, facilitating the initiation and continuation of inter-agent conversations.
The framework deploys a combination of an LLM-powered assistant agent and a user agent in an assistant-user cooperative scenario. The assistant agent, powered by LLMs (OpenAI, 2022), generates solutions and communicates these to the user agent. The user agent then executes the assistant’s recommendations, providing feedback to the assistant agent.
@@ 3.1.1. Role Specialization
3.1.2. Action Execution
After role specialization, assistant and user agents collaborate in an instruction-following manner to complete the assigned tasks. Human users initiate the action execution by inputting contract codes into the system. The system segments this code into a task queue, which is systematically processed through three distinct phases: Contract Analysis, Vulnerability Identification, and Comprehresive Report. The Contract Analysis phase involves preliminary assessments of the contract’s purpose and structure. During Vulnerability Identification phase, agents collaboratively identify and describe potential security weaknesses. Finally, the Comprehresive Report phase generates a comprehensive audit report.
@@ 3.1.2. Action Execution
Figure 2. Thought-Reasoning and Buffer-Reasoning Prompting Strategies
3.2. Task Queue
@@
3.2.1. Structured Prompting for Task Specialization
@@
Prompt engineering is a crucial element in achieving role specialization and efficient task execution within our framework. To enable this, we employ inception prompting (Li et al., 2023), a technique that defines the roles, tasks, and responsibilities of both assistant and user agents. Each inception prompt includes three essential components: the specified task prompt, assistant agent prompt, and user agent prompt. This structured approach transforms initial concepts into specific, actionable tasks that harness the unique capabilities of each agent. Building on this foundation, our framework incorporates two strategies: Thought-Reasoning and Buffer-Reasoning.
Thought-Reasoning prompt originates from the ReAct framework (Yao et al., 2022), which effectively combines reasoning and acting in language models to handle complex tasks with adaptive thought and action. Our approach builds on this foundation, extending ReAct’s synergy between reasoning and action to ensure that the model not only processes information sequentially but also actively interacts with external data sources as needed. Unlike single-query (Few-shot or CoT) to model, which may lead to surface-level analyses, this method continuously verifies and refines reasoning through targeted interactions, enhancing depth and accuracy in complex task-solving.
Figure 2 illustrates the Thought-Reasoning prompt template, adapted specifically for broad-spectrum vulnerability analysis. The specified task sets out the main goal—identifying various vulnerabilities within GPT’s knowledge base—by clearly defining expected outputs, including vulnerability types and detailed descriptions, alongside constraints to mitigate potential undesirable behaviors. The assistant agent and user agent provide structured guidance on roles and tasks, fostering collaboration between specialized agents to systematically uncover vulnerabilities in smart contracts.
Buffer-Reasoning prompt, derived from the foundational Buffer of Thoughts (BoT) (Yang et al., 2024) concept, enhances the model’s task comprehension by drawing upon high-level, context-specific thought templates, which is particularly crucial for complex domains. While ReAct is effective for many tasks, it often falls short in highly specialized areas, as LLMs generally lack domain-specific knowledge stores (Touvron et al., 2023). LLMs typically lack domain-specific knowledge vault, limiting their performance in specialized fields. Buffer-Reasoning prompt addresses this limitation by retrieving relevant thought templates from prior problem-solving processes, supporting in-context learning and bolstering the model’s understanding. Unlike ReAct, Buffer-Reasoning prompt inherently guides LLMs to engage in deep, step-by-step reasoning required for tackling complex tasks. Building on this foundational concept, LLM-SmartAudit utilizes an adapted Buffer-Reasoning prompt approach. This method combines thought-augmented reasoning with adaptive instantiation, prompting LLMs to integrate contextual information and systematically analyze the problem.
@@
Figure 2 presents an example of the specified task prompt for Buffer-Reasoning prompt, tailored for focused and detailed analysis of smart contracts. This prompt establishes a clear objective for the agents, concentrating on identifying Transactions Order Dependence (TOD) vulnerabilities. It directs agents to examine the contract’s logic and critical functions, particularly those involving fund transfers, resource allocation, or gas price manipulation, for TOD vulnerabilities. If a TOD vulnerability is identified, the assistant agent must provide a detailed description of the vulnerability and its potential impact. Conversely, if no TOD vulnerability is found, the agent outputs “NO TOD”.
3.2.2. Execution Mode